OpenStudy (anonymous):

1500 families with 2 children were selected randomly and the following data was recorded : no of girls : 0 , 1 , 2 no. of families : 211 , 814 , 475 if a family is chosen at random , find the probability that it has a) at most one girl b) at least one girl

OpenStudy (kropot72):

\[P(0\ girls)=\frac{211}{1500}\] \[P(1\ girl)=\frac{814}{1500}\] \[P(2\ girls)=\frac{475}{1500}\] \[P(at\ most\ 1\ girl)=P(0\ girls)+P(1\ girl)\] \[P(at\ least\ 1\ girl)=P(1\ girl)+P(2\ girls)\]

OpenStudy (anonymous):

what is the actually meaning of atmost

OpenStudy (kropot72):

At most 1 girl means no more than 1 girl. Therefore 'at most 1 girl' includes families with no girls and families with 1 girl.

OpenStudy (anonymous):

thank u

OpenStudy (kropot72):

You're welcome :)